OpenStudy (anonymous):

hey guys i need to complete the square for this problem: If 3x^2 – 2x + 7 = 0 Can someone help??:)

OpenStudy (anonymous):

sure but it is a pain

OpenStudy (anonymous):

ok thank you:))

OpenStudy (anonymous):

\[3x^2-2x=-7\] \[x^2-\frac{2}{3}x=-\frac{7}{3}\]

OpenStudy (anonymous):

that is to get the leading coefficient 1, otherwise you cannot complete the square.

OpenStudy (anonymous):

then take half of \[-\frac{2}{3}\] square it and add it to the other side

OpenStudy (anonymous):

\[(x-\frac{1}{3})^2=-\frac{7}{3}+\frac{1}{9}\]

OpenStudy (anonymous):

\[(x-\frac{1}{3})2=-\frac{20}{9}\]

OpenStudy (anonymous):

now take the square root. your answer is complex because you have a negative number on the right

OpenStudy (anonymous):

thank you so much!!! that is all i needed!!

OpenStudy (anonymous):

youre a great teacher:D

myininaya (myininaya):

satellite stop being so glamorous

OpenStudy (anonymous):

\[x-\frac{1}{3}=\pm\frac{20}{3}i\]

myininaya (myininaya):

satellite let me teach you splitting the middle please its easy

OpenStudy (anonymous):

hello my myininaya

OpenStudy (anonymous):

split the what???

OpenStudy (anonymous):

i factor by setting = 0 and solving. that is the only way i know how.

OpenStudy (anonymous):

oh forgot to complete the problem. \[x=\frac{1}{3}\pm\frac{\sqrt{20}}{3}i\]

OpenStudy (anonymous):

haha thank you, i'm studying for my cpt and this was confusing the heck out of me.